Find the largest value of parameter a such that equation 12sin2x * cos2x + (10a - 11)(cos4x - sin4x) = (2a-1)2 has at least one real solution.


1
Expert's answer
2021-10-04T18:41:51-0400

Solution:

"12\sin^2x \times \cos^2x + (10a - 11)(\cos^4x - \sin^4x) = (2a-1)^2
\\ \Rightarrow 12(1-\cos^2x) \times \cos^2x + (10a - 11)((\cos^2x)^2 - (\sin^2x)^2) = (2a-1)^2"

"\\ \Rightarrow 12(1-\cos^2x) \times \cos^2x + (10a - 11)((\cos^2x+\sin^2x)(\cos^2x - \sin^2x)) = (2a-1)^2"

"\\ \Rightarrow 12(1-\cos^2x) \times \cos^2x + (10a - 11)(1(2\cos^2x -1)) = (2a-1)^2"

"\\ \Rightarrow 12(1-\cos^2x) \cos^2x + (10a - 11)(2\cos^2x -1) = (2a-1)^2"

Put "\cos x=y"

"\\ 12(1-y^2) y^2 + (10a - 11)(2y^2 -1) = (2a-1)^2
\\ \Rightarrow12(1-y^2) y^2 + (10a - 11)(2y^2 -1) = (2a-1)^2
\\ \Rightarrow12y^2-12y^4 + 20ay^2-22y^2-10a+11 = 4a^2+1-4a"

"\\ \Rightarrow-12y^4 + 20ay^2-10y^2-6a+10-4a^2 = 0
\\ \Rightarrow6y^4 -10ay^2+5y^2+3a-5+2a^2 = 0
\\ \Rightarrow6y^4 +y^2(-10a+5)+3a-5+2a^2 = 0"

Put "y^2=t"

"6t^2 +t(-10a+5)+3a-5+2a^2 = 0
\\\Rightarrow6y^4 +y^2(-10a+5)+3a-5+2a^2 = 0"

For at least one real solution, discriminant "D\ge 0"

"D=B^2-4AC=(-10a+5)^2-4(6)(3a-5+2a^2)
\\=100a^2+25-100a-24(3a-5+2a^2)
\\=100a^2+25-100a-72a+120-48a^2
\\=52a^2-172a+145"

Now, "D \ge 0"

"52a^2-172a+145\ge0
\\\Rightarrow52\left(a-\frac{43}{26}\right)^2+\frac{36}{13}\ge \:0
\\ \Rightarrow 52\left(a-\frac{43}{26}\right)^2\ge \:-\frac{36}{13}
\\ \Rightarrow \left(a-\frac{43}{26}\right)^2\ge \:-\frac{9}{169}"

"\mathrm{True\:for\:all}\:a"

No largest value of 'a' is defined.